# Defining implication relation for classical logic

### v11

@@ 7 $\phi(a,b)$ 3: Universal instantiation
8 $\neg\phi(a,b)$ 6: Universal instantiation
9 $\bot$ 7,8: LNC
@@
  • $\forall x\in X=\neg(\exists x\in X)\neg$, or equivalently, $\forall x\in X(\phi)\Leftrightarrow\forall x(\neg\chi(x)\vee\phi)$.
@@
The universal bounded quantifier $\forall x\in X$ can not be defined by $\forall x\in X(\phi)\Leftrightarrow\forall x(\chi(x)\to\phi)$ since $\chi(x)\to\phi$ and $\neg\chi(x)\vee\phi$ are not equivalent – however, if they are theorems (they are theorems if one of them is a theorem), then they are equivalent.
@@ 9 Concluding remarks
According to the original meaning of implication, the formula $\neg\phi\vee\psi\to(\phi\to\psi)$ is not a tautology in general, so that the equivalence $(\phi\to\psi)\leftrightarrow\neg\phi\vee\psi$, on which the material implication of classical logic is based, is generally invalid. Defining implication in accordance with its original meaning is the key to avoid invalid results in logic. The logic IRL presented in this work has the expected property that the invalid theorem $\neg\phi\vee\psi\to(\phi\to\psi)$ of classical logic is removed while fundamental laws in classical logic, such as LEM and double negation, LNC and ECQ, conjunction elimination and disjunction introduction, and transitivity of implication and disjunctive syllogism, are all retained in IRL.
Based on information provided with IRL, more points are summarized as follows.
@@
    Truth value of implication. Since implication is not truth functional, valuation of $\phi\to\psi$ can not be generally done by valuation of $\phi$ and $\psi$, instead, it can be “estimated” via the procedure: (a) $\phi\to\psi$ is a tautology, if and only if $\neg\phi\vee\psi$ is a tautology; (b) $\phi\to\psi$ is a contradiction, if $\neg\phi\vee\psi$ is a contradiction (i.e. $\phi$ is a tautology and $\psi$ is a contradiction); (c) the T-set of $\phi\to\psi$ is a subset (possibly empty) of the T-set of $\neg\phi\vee\psi$, if $\neg\phi\vee\psi$ is a contingency. T-sets of primitive implications, like any primitive propositions, are given (e.g. by physical mechanisms) rather than derived.
@@
    Classical logic is sound and complete with regard to its own semantics. In the sense that it contains theorems that are not generally valid with regard to semantics based on the original meaning of implication, classical logic is less “sound” than “complete” (actually it is post-complete), while some weaker systems are in the opposite.
@@
    The logic IRL of this work, in which the generally invalid classical theorems – including especially the class of $\neg\phi\vee\psi\to(\phi\to\psi)$ – are removed, should be used instead of classical logic for the general case (the system contains any types of propositions including especially contingencies). IRL and classical logic are the same if and only if there is no contingent primitive propositions in a system considered.
@@
    When an interpretation or model is given for a formal system, all statements of that system are not contingent (actually, an interpreted statement is either true or false, irrelevant to “tautology, contradiction or contingency”). It is common that a formal theory has an intended interpretation or standard model or usual model. For example, Tarski’s axioms (for Euclidean geometry) has its standard model of Euclidean plane geometry, and the standard model of Peano axioms is natural numbers. In such situations where there are no contingent primitive sentences, classical logic can be used safely.

### v12

@@ 6     $\phi(a,b)$ 3: Universal instantiation
7     $\neg\phi(a,b)$ 5: Universal instantiation
8     $\bot$ 6,7: LNC
9 $\exists x\forall y(\phi(x,y))\wedge\neg\forall y\exists x(\phi(x,y))\to\bot$ 1–8: Conditional proof
10 $\exists x\forall y(\phi(x,y))\to\forall y\exists x(\phi(x,y))$ 9: Implication and inconsistency
Meta Rule 8.0.1 (Equivalence of implication theorem and disjunction theorem in FOL).
$(\vdash\forall x\,(\phi(x)\to\psi(x)))\Leftrightarrow(\vdash\forall x\,(\neg\phi(x)\vee\psi(x))).$
Proof.
1 $\vdash\forall x\,(\phi(x)\to\psi(x))$ Premise
2     $\forall x\,(\phi(x)\to\psi(x))$ 1 (assumed theorem)
3     $\phi(y)\to\psi(y)$ 2: Universal instantiation ($y$ is arbitrary)
4 $\vdash\phi(y)\to\psi(y)$ 2–3: Unconditional proof
5 $\vdash\neg\phi(y)\vee\psi(y)$ 4: Meta Rule 3.4.1 of IRL
6     $\neg\phi(y)\vee\psi(y)$ 5 (theorem indicated in an outer proof)
7     $\forall x\,(\neg\phi(x)\vee\psi(x))$ 6: Universal generalization (as $y$ is arbitrary)
8 $\vdash\forall x\,(\neg\phi(x)\vee\psi(x))$ 6–7: Unconditional proof
The converse can be proved in the reverse order. ∎
Bounded quantifiers
@@ 8 $\vdash\forall x\,(\neg\phi(x)\vee\psi(x))$ 6–7: Unconditional proof
  • $\forall x\in X\,(\phi)=\neg\exists x\in X\,(\neg\phi)\Leftrightarrow\neg\exists x\,(\chi(x)\wedge\neg\phi)\Leftrightarrow\forall x\,\neg(\chi(x)\wedge\neg\phi)$, i.e., $\forall x\in X\,(\phi)\Leftrightarrow\forall x\,(\neg\chi(x)\vee\phi)$.
@@
The universal bounded quantifier $\forall x\in X$ can not be defined by $\forall x\in X\,(\phi)\Leftrightarrow\forall x\,(\chi(x)\to\phi)$ since $\chi(x)\to\phi$ and $\neg\chi(x)\vee\phi$ are not equivalent in IRL – however, if $\forall x\in X\,(\phi)$ represents a theorem, the two expressions are equivalent according to the above Meta Rule 8.0.1.
9 Discussion
Normal situation
An arbitrary proposition can be tautological, contradictory, or contingent. Most generally, there are possibly infinitely many T-set values for a contingency ($10<x<$ 1) whereas there is only one value for a tautology ($x=1$) and one value for a contradiction ($x=0$). This is just similar to the unit interval $[0,1]$ (no matter it is different as totally ordered) where the number of interior points is uncountable while the number of boundary points is only two. In this sense, it is the normal situation for a proposition being contingent rather than tautological or contradictory. Subsequently in this section, when saying “in the normal situation” we mean propositions mentioned are all contingent.
Interpretation of Generalized ECQ of implication
In IRL, the Generalized ECQ $\psi\to(\phi\to\psi)$ does not hold while the Generalized ECQ of implication $(\phi\to\psi)\to(\chi\to(\phi\to\psi))$ does.
Generalized ECQ $\psi\to(\phi\to\psi)$ means that, for any propositions $\phi$ and $\psi$, there is always an implication from $\phi$ to $\psi$ in the normal situation (where $\psi$ can be true so $\phi\to\psi$ can be true). This is not consistent with the original meaning of implication as it is not ensured for a “certain mechanism” of implication in this situation.
Interpretation of ECQ
A simple conditional proof of ECQ $\phi\wedge\neg\phi\to\psi$:
1 $\phi$ Premise 2 $\neg\phi$ Premise 3 $\phi\vee\psi$ 1: Disjunction introduction 4 $\psi$ 2,3: Disjunctive syllogism
Hence ECQ is necessary for disjunction introduction and disjunctive syllogism which do not conflict to the original meaning of implication. Therefore, ECQ contributes to the goal of strength for the system (cf. Subsection 2.6).
Furthermore, ECQ $\bot\to\phi$ and its dual ϕitalic-$\phi\to\top$ can be interpreted to intuition as “if an always-false thing is (was) true, then anything is (would be) true” and “in any cases, an always-true thing is true”, respectively, as similarly mentioned, e.g., in (Ceniza, 1988).
ECQ is an abnormal situation, for which a “standard (set of rules for) interpretation” may be explored for practical purpose. For example, given a conditional $P\to Q$: if $P$ is a contradiction (e.g. it is well known that $P$ is false), then $P\to Q$ can be interpreted as “if the false thing like $P$ were true, anything including $Q$ should be true”; if $Q$ is a tautology (e.g. it is a well known fact), then $P\to Q$ can be interpreted as “in any cases including $P$, an always true thing like $Q$ is true”.
@@ 10 Concluding remarks
By the original meaning of implication, the formula $\neg\phi\vee\psi\to(\phi\to\psi)$ is generally invalid, so the material implication $(\phi\to\psi)\leftrightarrow\neg\phi\vee\psi$ is not valid in general. Defining implication according to its original meaning is the key to avoid invalid results. The logic IRL of this work has the expected property that the invalid class of $\neg\phi\vee\psi\to(\phi\to\psi)$ is removed while fundamental laws like LEM and double negation, LNC and ECQ, conjunction elimination and disjunction introduction, and transitivity of implication and disjunctive syllogism, are all retained. Based on information provided with IRL, more points are summarized as follows.
@@
    Truth value of implication. Since implication is not truth functional, valuation of $\phi\to\psi$ can not be generally done by valuation of $\phi$ and $\psi$, but it can be “estimated” via the procedure: (a) $\phi\to\psi$ is a tautology, if and only if $\neg\phi\vee\psi$ is a tautology; (b) $\phi\to\psi$ is a contradiction, if $\neg\phi\vee\psi$ is a contradiction (i.e. $\phi$ is a tautology and $\psi$ is a contradiction); (c) the T-set of $\phi\to\psi$ is a subset (possibly empty) of the T-set of $\neg\phi\vee\psi$, if $\neg\phi\vee\psi$ is a contingency. T-sets of primitive implications, like any primitive propositions, are given (e.g. by physical mechanisms) rather than derived.
@@
    Classical logic is sound and complete with regard to its own semantics. In the sense that it contains theorems that are not generally valid with respect to semantics based on the original meaning of implication, classical logic is less “sound” than “complete” (actually it is post-complete), while some weaker systems are in the opposite.
@@
    The logic IRL of this work, in which the generally invalid classical theorems, including especially the class of $\neg\phi\vee\psi\to(\phi\to\psi)$, are removed, should be used instead of classical logic for the general case (the system contains any types of propositions including especially contingencies). IRL and classical logic are the same if and only if there is no (primitive) contingent propositions in a system.
@@
    An interpreted formal system has no contingent statements. Precisely, for a system under a given interpretation, every sentence is simply either true or false, no longer relevant to “tautological, contradictory or contingent”. It is common that a theory has an intended interpretation or a standard model or usual model. For example, Tarski’s axioms for Euclidean geometry has its standard model of Euclidean plane geometry, and the standard model for Peano axioms and Robinson arithmetic is the natural numbers. In such contexts when a system contains no contingencies, classical logic can be used safely.